shamrock has 900 pounds of milk that is 2% butterfat, how much butterfat do they have to add to raise the butterfat content to 8%?
Answer by mananth(12270)

percent ---------------- quantity
Milk 2 ---------------- 900 pounds
Butter fat 100 ---------------- x pounds
Mixture 8 ---------------- 900 + x pounds

sum of individual components = quantity in mixture
2*900+100x=8(900+x)
1800+100x=7200 +8x
100x-8x =5400
92x=5400
x= 5400 / 92
x= 58.7
58.7 pounds of Butter fat has to be added